# Transforming Norm-based To Graph-based Spatial Representation for Spatio-Temporal Epidemiological Models

### 22 Feb 2024 v1

@@ Abstract
Pandemics, with their profound societal and economic impacts, pose significant threats to global health, mortality rates, economic stability, and political landscapes. In response to the persistent challenges posed by emerging and reemerging pandemics, numerous studies have employed spatio-temporal models to enhance our understanding and management of these complex phenomena. These spatio-temporal models can be roughly divided into two main spatial categories: norm-based and graph-based trade-offering between accuracy, computational burden, and representational feasibility. In this study, we explore the ability to transform from norm-based to graph-based spatial representation for these models. We introduce a novel framework for this task together with twelve possible implementations using a wide range of heuristic optimization approaches. Our findings show that by leveraging agent-based simulations and heuristic algorithms for the graph node’s location and population’s spatial walk dynamics approximation one can use graph-based spatial representation without losing much of the model’s accuracy and expressiveness. For three real-world cases, the best-performing algorithmic configuration archives 94% accuracy presence, on average. Moreover, an analysis of synthetic cases shows the proposed framework is relatively robust, as fluctuation in both spatial and temporal dynamics is not badly reflected by the framework’s performance.
@@ 1 Introduction
The historical trajectory of pandemics reveals a profound impact on societies across centuries, leaving an indelible mark on mortality rates, economic stability, and political landscapes [1]. In the last several decades, the persistent threat of emerging and reemerging pandemics and epidemics looms large over humanity [2]. Indeed, recent outbreaks such as the severe acute respiratory syndrome (SARS) in 2003, the H1N1 influenza pandemic in 2009, the Middle East respiratory syndrome coronavirus (MERS-CoV) in 2012, the Ebola virus in 2014, and the ongoing global severe acute respiratory syndrome coronavirus 2 (SARS-CoV-2) pandemic, underscore the urgency of comprehending, preparing, and managing pandemics [3, 4, 5, 6, 7].
To this end, multiple mathematical and computational models have been proposed to ameliorate our understanding of the pandemic spread and advance our ability to control it [8, 9, 10, 11]. These models can be roughly divided into two primary groups: statistical models and mechanistic models. The former group employs a data-driven approach without making explicit assumptions about the underlying dynamics, while the latter relies on theoretical principles to elucidate the mechanisms governing pandemic dynamics. Statistical models commonly leverage statistical and machine learning (ML) methods to project a diverse array of outcomes. Within this category, a multitude of models has been developed utilizing various techniques, including auto-regressive time series methods, Bayesian optimization techniques, and, more recently, deep learning models [12, 13, 14, 15, 16]. These models have been shown to provide promising results on the pandemic they trained on but they usually do not generalize well on other cases, require large amounts of data to obtain reasonable performance, and are very sensitive to noise. The latter group, exemplified by the SIR model, operates based on a predefined set of rules or processes governing interactions among individuals within a population [17, 18, 19, 20]. At the forefront of this approach stands the Susceptible-Infected-Recovered (SIR) model [21, 22, 23], originally proposed by Kermack and McKendrick [24]. Models belonging to this group require more effort from the modeler as one needs to manually detect, define, and formalize the main interactions and processes occurring in a pandemic for the model to capture the dynamics correctly. That said, once done efficiently, models from this group are more generalized, adaptable, explainable, and require less data.
@@
One can easily notice that the norm-based representation is more expressive than the graph-based one. Nonetheless, this expressiveness can be computationally overwhelming while relatively easy to model and comprehend compared to the more abstract graph-based approach. Hence, in this study, we proposed a novel algorithm that utilizes an agent-based simulation (ABS) [42, 43] representation of a norm-based spatio-temporal model to obtain a numerically similar graph-based spatio-temporal model. We evaluate the proposed model on three different temporal epidemiological models as well as three different pandemic settings to explore the algorithm’s performance.

### 26 Feb 2024 v2

@@ Abstract
Pandemics, with their profound societal and economic impacts, pose significant threats to global health, mortality rates, economic stability, and political landscapes. In response to these challenges, numerous studies have employed spatio-temporal models to enhance our understanding and management of these complex phenomena. These spatio-temporal models can be roughly divided into two main spatial categories: norm-based and graph-based. Norm-based models are usually more accurate and easier to model but are more computationally intensive and require more data to fit. On the other hand, graph-based models are less accurate and harder to model but are less computationally intensive and require fewer data to fit. As such, ideally, one would like to use a graph-based model while preserving the representation accuracy obtained by the norm-based model. In this study, we explore the ability to transform from norm-based to graph-based spatial representation for these models. We first show no analytical mapping between the two exists, requiring one to use approximation numerical methods instead. We introduce a novel framework for this task together with twelve possible implementations using a wide range of heuristic optimization approaches. Our findings show that by leveraging agent-based simulations and heuristic algorithms for the graph node’s location and population’s spatial walk dynamics approximation one can use graph-based spatial representation without losing much of the model’s accuracy and expressiveness. We investigate our framework for three real-world cases, achieving 94% accuracy preservation, on average. Moreover, an analysis of synthetic cases shows the proposed framework is relatively robust for changes in both spatial and temporal properties.
@@ 1 Introduction
Pandemics over history caused intense impact on societies across centuries, leaving an indelible mark on mortality rates, economic stability, and political landscapes [1]. In the last decades, the persistent threat of emerging and reemerging pandemics and epidemics looms large over humanity [2]. Indeed, recent outbreaks such as the severe acute respiratory syndrome (SARS) in 2003, the H1N1 influenza pandemic in 2009, the Middle East respiratory syndrome coronavirus (MERS-CoV) in 2012, the Ebola virus in 2014, and the ongoing global severe acute respiratory syndrome coronavirus 2 (SARS-CoV-2) pandemic, underscore the urgency of comprehending, preparing, and managing pandemics [3, 4, 5, 6, 7].
To this end, multiple mathematical and computational models have been proposed to ameliorate our understanding of the pandemic spread and advance our ability to control it [8, 9, 10, 11]. These models can be roughly divided into two primary groups: statistical models and mechanistic models. The former group employs a data-driven approach without making explicit assumptions about the underlying dynamics, while the latter relies on theoretical principles to elucidate the mechanisms governing pandemic dynamics. Statistical models, on the other hand, commonly leverage statistical and machine learning (ML) methods to project a diverse array of outcomes. Within this category, a multitude of models has been developed utilizing various techniques, including auto-regressive time series methods, Bayesian optimization techniques, and, more recently, deep learning models [12, 13, 14, 15, 16]. These models have been shown to provide promising results on the pandemic they trained on but they usually do not generalize well on other cases, require large amounts of data to obtain reasonable performance, and are very sensitive to noise. The latter group, exemplified by the SIR model, operates based on a predefined set of rules or processes governing interactions among individuals within a population [17, 18, 19, 20]. At the forefront of this approach stands the Susceptible-Infected-Recovered (SIR) model [21, 22, 23], originally proposed by Kermack and McKendrick [24]. Models belonging to this group require more effort from the modeler as one needs to manually detect, define, and formalize the main interactions and processes occurring in a pandemic for the model to capture the dynamics correctly. That said, once done efficiently, models from this group are more generalizable, adaptable, explainable, and require less data.
@@
One can easily notice that the norm-based representation is more expressive than the graph-based one. Nonetheless, this expressiveness can be computationally overwhelming while relatively easy to model and comprehend compared to the more abstract graph-based approach. Therefore, it is interesting to pose the question of whether we can transform models from norm-based to graph-based representations, in a way that preserves the model’s accuracy in representing the epidemiological dynamics. Hence, in this study, we proposed a novel algorithm that utilizes an agent-based simulation (ABS) [42, 43] representation of a norm-based spatio-temporal model to obtain a numerically similar graph-based spatio-temporal model. We evaluate the proposed model on three different temporal epidemiological models as well as three different pandemic settings to explore the algorithm’s performance.
@@ 5 Discussion
Acknowledgment
The author wishes to thank Sveta Hardak-Nissan for teaching me the professional maturity required for fundamental research.